Solve the following ordinary differential equations:

d²y/dx²+dy/dx+y=0


1
Expert's answer
2021-01-04T17:16:11-0500

For solving equation : "y^{(2)}+y^{(1)}=-y" .

The solution is "y=e^{q*x}" .

By substituting such expression into the differential equation can be formulated as the equation on "q" : "(q^{2} + q - 1)*e^{q*x}=0" .

Where we substitute: "y^{(2)} = q^{2}*y" and "y^{(1)} = q" .

"e^{q*x} \neq 0" as the exponential function will always be non-zero  for any finite "x" so zeros must come from:

"q^2 + q + 1 =0"

solution of obtained equation is

"q_{+,-} = -\frac{1}{2} \pm \frac{i 3^{1/2}}{2}"

so solutions of the equation are

"y_{+,-} = C_{+,-}* exp(-\frac{1}{2} \pm \frac{i 3^{1/2}}{2})"

And the general solution of equation is

"y = C_{+}* exp(-\\frac{1}{2} + \\frac{i 3^{1\/2}}{2}) + \ufeffC_{-}* exp(-\\frac{1}{2} - \\frac{i 3^{1\/2}}{2})"

also the general solution can be presented in the trigonometrical form, by substituting the formula into Euler's identity "e^{(a+i*b)}=e^a(\cos{b}+i*\sin{b})" the general solution will be given by

"y = C_{1}* e^{-\\frac{x}{2}}*\\sin{ \\frac{\\sqrt{3}}{2}x} + \ufeff C_{2}* e^{-\\frac{x}{2}}*\\cos{ \\frac{\\sqrt{3}}{2}x}"